# Seeing Clearly, Answering Incorrectly: A Multimodal Robustness Benchmark for Evaluating MLLMs on Leading Questions

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have exhibited impressive capabilities in visual understanding and reasoning, providing sightly reasonable answers, such as image descriptions. This has spurred extensive research on the evaluation of MLLMs. Most evaluation benchmarks assume that incorrect answers indicate a lack of understanding of the visual content. However, our findings reveal that, in many cases, MLLMs answer questions incorrectly despite correctly understanding the visual content. This suggests that incorrect answers do not necessarily imply a lack of comprehension but may instead result from lacking robustness to leading questions. To comprehensively measure MLLMs’ understanding capability and robustness to leading questions, we introduce a MultiModal Robustness benchmark (MMR). MMR contains paired positive and negative questions across 12 categories, meticulously annotated by humans. We evaluate 18 leading MLLMs on the MMB benchmark, revealing that MLLMs suffer from fragility to leading questions despite understanding the visual content. To enhance MLLMs’ understanding capability and robustness, we further present a training set with paired positive and negative visual question-answer samples. Experiments verify that MLLMs’ robustness can be significantly enhanced by tuning on this new training set. The benchmark, training set, and code can be found at https://github.com/BAAI-DCAI/Multimodal-Robustness-Benchmark.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have garnered significant attention for their impressive capabilities in tackling diverse multimodal tasks, making notable strides in visual comprehension and reasoning [1; 2; 3]. This has spurred the development of various evaluation methods to assess the capabilities of MLLMs across different dimensions. These evaluation methods primarily encompass general benchmarks and hallucination benchmarks. General evaluation frameworks, such as MMBench [4], MMMU [5], VQA-v2 [6], GQA [7], ScienceQA [8], and et al., scrutinize models for their proficiency in generic and knowledge-based question and answer (VQA) tasks. In contrast, hallucination evaluation methods focus on evaluating the discrepancy between generated responses and the visual content. These methods typically add interference to the text (e.g., object [9; 10; 11; 12; 13; 14; 15; 16; 17], knowledge [9; 14; 15; 17], relation [18; 19; 15; 16], attribute [18; 20; 13; 14; 15; 16; 17] interference) to see if the MLLM can generate responses that are consistent with the visual content.
Figure 1: Left: Examples show that MLLMs understand content but answer incorrectly. Right: Performance comparison on positive and negative questions for the same visual content.
Most MLLM evaluation methods assume that incorrect answers indicate a lack of understanding of the visual content. However, we unveil a critical and overlooked phenomenon: MLLMs may produce incorrect answers despite correctly comprehending the visual content (as shown in Fig. 1). This suggests that MLLMs are susceptible to subtle interference on the question, hindering their ability to leverage their knowledge effectively. However, existing benchmarks and evaluation metrics, primarily focused on answer accuracy, fail to assess this phenomenon adequately. In this paper, we aim to construct a new benchmark for 1) assessing whether the model understands the visual content and 2) evaluating the robustness of MLLMs to leading questions.
To this end, we introduce the MultiModal Robustness benchmark (MMR). Each sample is manually labeled with one of 12 types of questions, with paired positive and leading questions. We design multiple-choice questions with answers and intrusive explanations to prevent the model from guessing. To quantify the evaluation, we devise a novel metric to measure the likelihood of a model’s failure when induced by a leading question, despite its comprehension of the visual content.
We evaluate 18 leading MLLMs on the MMR benchmark, as shown in Tab. 2. Surprisingly, these MLLMs perform unsatisfactorily in terms of leading questions. We attribute this issue to the scarcity of robustness training data. Previous datasets primarily consist of positive questions, with only a few including negative samples (such as LRV [9]) aiming at mitigating object hallucination. However, most existing training data is insufficient to effectively enhance the robustness of MLLMs, as shown in Tab. 3. To address this challenge, we introduce a data construction pipeline. We first extract visual information from images, including text/numbers, objects, attributes, relationships, and local/global comprehension. Based on this information, we generate paired positive and negative samples.
In summary, the contributions are three folds: 1) We introduce a human-annotated multimodal robustness benchmark (MMR) and new evaluation metrics for measuring MLLMs’ understanding capability and robustness to leading questions. 2) We present a data construction pipeline to enhance the understanding ability and robustness by generating paired positive and negative samples. 3) We extensively experiment on the MMR and general benchmarks. Results show that current MLLMs are severely affected by leading questions, which our proposed dataset can effectively mitigate.
@@ 2 Related Works
Multimodal large language models (MLLMs).
@@
Building on the success of Large Language Models, recent research has increasingly focused on MLLMs [21; 1]. This shift aims to achieve better cross-modality understanding and reasoning. Currently, a significant amount of research is exploring various aspects of MLLMs, including the design of structures [22; 23], data construction [24; 25; 26], training strategies [27; 28; 29], and lightweight MLLMs [3; 30]. For a comprehensive review of these topics, readers are encouraged to investigate surveys [31; 32; 33] dedicated to several areas.
Evaluations of MLLMs.
@@
The assessment of the MLLM’s capabilities is mainly divided into two main categories: general capability [8] and hallucination benchmarks [34]. General capability benchmarks can be divided into single-task benchmarks and comprehensive benchmarks. Single-task benchmarks, such as VQA [6], OK-VQA [35], VQA-v2 [6], GQA [7], and MS-COCO [36], are designed to assess the performance of models on a specific defined task. Comprehensive benchmarks, such as MMBench [4], SEED [37], MMMU [5], and MMStar [38] aim to assess MLLMs’ general multi-modal perception (such as color, position, and celebrity recognition.) and reasoning capabilities (such as numerical calculation and knowledge-based reasoning). Hallucination benchmarks mainly focus on evaluating the discrepancy between text responses generated by MLLMs and provided visual content [34]. Most hallucination benchmarks focus on object [39; 9; 10; 11; 12; 13; 14; 15; 16; 17], knowledge [9; 14; 15; 17], relation [18; 19; 15; 16], attribute [18; 20; 13; 14; 15; 16; 17], spurious images [40; 41] hallucinations. However, current benchmarks assume correct answers indicate understanding, while incorrect answers do not. We find that MLLMs answer incorrectly despite understanding the visual content.
Multimodal instruction-tuning data generation.
The generation of synthetic data for multimodal instruction tuning can be divided into two main methods: text-only-based and image-based generation. Text-only methods include LLaVA [22], which uses the text capabilities of GPT-4 [42] to build a multimodal dataset, InstructBLIP [43], which integrates VQA datasets focused on academic tasks to enhance visual comprehension, and LRV [9], which generates positive and negative samples using image bounding boxes for robust tuning. Image-based methods employ multimodal large language models to generate caption data [44; 24], VQA [7; 8; 45] complex reasoning instruction data [46; 47; 3; 48] and counterfactual data to address specific scenarios [18]. SPHINX-X [49] assembles an OCR-focused dataset from large-scale PDFs, converting text annotations and bounding boxes into a unified question-answering format.
3 MMR Benchmark
Benchmark Design.
As shown in Fig. 2, we design paired positive and negative questions for each sample. Each question has four brief options, with only one correct answer. The positive questions assess the model’s ability to understand the visual content correctly, while the negative questions evaluate the model’s robustness in answering correctly despite misleading prompts.
To achieve a comprehensive evaluation, we manually construct 300 positive and 300 leading negative questions across three levels: character, attribute, and context. Character-level questions prompt identifying elements like characters or numbers, while attribute-level questions focus on properties such as color, texture, and quantity. Context-level inquiries delve into higher-level concepts like emotions, culture, and common sense. The positive questions aim to evaluate the model’s understanding ability, while the misleading ones challenge its resistance to interference. For instance, character-level misleading questions alter elements like characters or numbers, attribute-level ones confuse properties, and context-level prompts introduce complex concepts, thoroughly testing the model’s resilience against misleading information.
@@ 3 MMR Benchmark
Figure 2: Examples in the MMR benchmark. The colors    ,    ,    , denote character, attribute, and context-level, respectively. POS and NEG denote the positive and negative questions, ANS
denotes the answer.
Evaluation metrics.
We categorize the results into four groups: 1) Understanding and Robustness (UR), 2) Understanding but Fragility (UF), 3) not Understanding and Rigorous (NR), and 4) not Understanding and Fragility (NF). To quantify the effect of negative questions in misleading the model despite its comprehension of visual content, we introduce the “Misleading Rate (MR)”. Additionally, we introduce the “Robustness Accuracy (RA)” to evaluate the actual understanding capability of MLLMs. The formulation is as follows:
$\text{MR}=\frac{N_{UF}}{N_{UR}+N_{UF}}$ (1)
$\text{RA}=\frac{N_{UR}}{N_{UR}+N_{UF}+N_{NR}+N_{NF}}$ (2)
Where $N_{\text{i}}$ ($\text{i}\in\{UR,UF,NR,NF\}$) represents the number of samples. By combining these two evaluation metrics, our benchmark can not only reflect the model’s basic understanding capabilities but also reveal its robustness to misleading prompts.
4 MMR-data Construction
Table 1: Comparison between different instruction tuning datasets in MLLMs.
Characteristics LLaVA ALLaVA LRV Ours
Hard negative instructions?
Paired positive-negative data?
Information extraction?
Average conversation rounds 2.35 1 1 6.82
Figure 3: Overview of data generation. 1) Key information within the image is extracted using GPT-4V. 2) The extracted key information is then used to generate paired positive and negative samples. 3) Filtering low-quality samples.
To enhance MLLMs’ understanding capability and robustness, we propose a data construction pipeline that utilizes GPT-4V to generate paired positive and negative samples for instruction tuning, as shown in Fig. 3. The distinctions between the proposed dataset and previous conversation datasets are illustrated in Tab. 1.
Information extraction. Previous approaches that generate instruction-tuning data for multimodal conversations with GPT-4V primarily fall into two categories: direct generation and annotation-driven generation. Direct generation methods, such as ALLaVA [46], rely entirely on the basic ability of GPT-4V to synthesize samples. In contrast, annotation-driven generation methods, such as LRV [9], utilize existing image annotations (e.g.object bounding boxes and textual descriptions) as additional information for text-only GPT-4 to generate conversational samples. However, these methods fail to fully utilize the fine-grained details. In contrast, our data construction pipeline adopts a different approach by implicitly and comprehensively extracting this detailed information directly from the image. This includes 1) text or numbers (if present), 2) objects and people, 3) object attributes (e.g., colors, textures, locations) and human characteristics (e.g., postures, dresses), 4) interrelationships between people and objects, 5) events or activities, and 6) the overall feeling and understanding evoked by the image. The detailed prompts are in the Appendix.
Instruction tuning data generation. The most relevant work to our data generation approach is LRV [9], which uses the Visual Genome [50] dataset to guide GPT-4 in generating positive and negative samples about object presence and manipulation knowledge. However, LRV has two significant limitations: limited fine-grained information and unpaired negative samples. It relies solely on bounding boxes and object labels, omitting textual descriptions, numerical attributes, and deeper image comprehension. Additionally, LRV’s negative samples are not explicitly paired with positive counterparts, reducing training effectiveness. To address these challenges, we generate positive samples using extracted information and construct negative samples that directly contradict the positive ones. This ensures a strong contrast for more effective model training, providing richer context and paired positive and negative samples. Detailed prompts are in the Appendix.
Sample filtering. We collect part of COCO [51] images from the LLaVA 158k and construct paired positive and negative samples. We filter the samples by keyword matching, removing conversations with uncertain answers (e.g., “uncertain”) and redundant phrases (e.g., “in the image”).
5 Visual Instruction Tuning
Framework.
Training.
We employ a two-stage training strategy. In the first stage, the visual and language alignment stage, we use Bunny-pretrain-LAION-2M [3] data to train only the connector while freezing the vision encoder and the language model. In the second stage, we fine-tune both the connector and the language model using the generated MMR dataset (MMR-data). To further enrich the dataset’s diversity, we include OCR data (OCR-VQA [54]), visual question-answering data (VQA-v2 [6], GQA [7], OK-VQA [55], A-OKVQA [35], Visual Genome [50], and RefCOCO [56]). Additionally, to maintain the language capabilities of the MLLM, we incorporated text-only data from WizardLM [57].
@@ 5 Visual Instruction Tuning
Table 2: Comparison to state-of-the-art MLLMs on the MMR benchmark. §We evaluate the officially released checkpoint by ourselves. Abbreviations: Char/Num. (Character/Number), Pres. (Presence), Color/Tex. (Color/Texture), Num. (Number), Shape (Shape), Posture (Posture), Pos. (Position), Abstract. (Abstract Knowledge), Concrete. (Concrete Knowledge), Expert (Expertise), Act. (Activity), Rel. (Relationships). $\uparrow$: higher is better, $\downarrow$: lower is better.
@@
GPT4o [42] 72.50 68.18 66.67 45.83 87.5 70.83 50.00 68.18 76.19 70.97 83.33 63.64 69.00
Gemini-pro-vision [1] 42.50 50.00 41.67 25.00 83.33 50.00 45.45 40.91 47.62 45.16 70.83 45.45 48.67
Qwen-VL-max [58] 67.50 72.73 66.67 41.67 79.17 62.5 63.64 77.27 80.95 61.29 79.17 72.73 68.33
Claude3-Opus-V [59] 35.00 22.73 12.50 16.67 33.33 16.67 22.73 45.45 33.33 25.81 37.50 40.91 28.67
Step-1V [60] 60.00 54.55 58.33 20.83 70.83 54.17 31.82 54.55 57.14 45.16 79.17 50.00 53.33
Glm-4V [61] 60.00 54.55 54.17 29.17 58.33 41.67 27.27 72.73 47.62 35.48 70.83 45.45 50.00
@@
LLaVA-1.6-34B [62] 75.00 68.18 66.67 41.67 79.17 54.17 72.72 81.81 71.42 64.52 79.17 68.18 68.67
Mini-Gemini-HD-34B [23] 62.50 63.64 70.83 54.17 79.17 62.50 72.73 86.36 85.71 54.84 19.17 68.18 69.00
Yi-VL-34B [63] 52.50 63.64 70.83 41.67 75.00 37.50 59.09 68.18 57.14 48.39 70.83 63.64 58.33
InternVL-Chat-V1-5 (26B) [64] 62.5 59.09 66.67 41.67 66.67 41.67 54.55 63.64 66.67 45.16 79.17 72.73 59.67
MiniCPM-Llama3-V [65] 37.5 45.45 50.00 16.67 41.67 37.5 36.36 68.18 33.33 29.03 41.67 54.55 40.33
Cogvlm2-llama3 [66] 60.00 63.64 54.17 37.5 70.83 33.33 40.91 50.00 85.71 41.94 62.50 50.00 54.00
LLaVA-RLHF (7B) [68] 7.50 36.36 33.33 33.33 50.00 16.67 9.09 59.09 38.10 22.58 50.00 31.82 30.67
Deepseek-VL-7B-Chat [29] 52.50 54.55 54.17 37.5 62.5 25.00 18.18 54.55 52.38 35.48 75.00 50.00 47.67
Mplug-owl2-llama2-7B [69] 32.50 63.64 58.33 20.83 62.50 37.50 13.64 54.55 47.62 25.81 58.33 31.82 41.33
Bunny-MMR-3B 60.0 59.09 58.33 25.0 83.33 50.0 54.55 68.18 57.14 51.61 79.17 54.55 58.33
@@
GPT4o [42] 9.38 16.67 23.81 26.67 4.55 19.05 38.89 28.57 15.79 24.14 13.04 22.22 19.46
Gemini-pro-vision [1] 29.17 31.25 41.18 45.45 13.04 40.00 33.33 52.63 44.44 48.15 19.05 23.08 34.82
Qwen-VL-max [58] 22.86 11.11 23.81 28.57 5.00 25.00 30.00 19.05 19.05 29.63 9.52 15.79 20.23
Claude3-Opus-V [59] 26.32 50.00 72.73 55.56 50.00 71.43 50.00 37.5 46.15 60.00 43.75 30.77 48.50
Step-1V [60] 14.29 25.00 26.32 61.54 5.56 40.91 61.11 33.33 33.33 44.00 9.52 21.43 30.43
Glm-4V [61] 27.27 36.84 35.00 56.25 33.33 52.38 53.85 20.00 47.37 57.69 19.05 37.50 38.78

### v2

# Unveiling the Ignorance of MLLMs: Seeing Clearly, Answering Incorrectly

@@ Abstract
Multimodal Large Language Models (MLLMs) have displayed remarkable performance in multi-modal tasks, particularly in visual comprehension. However, we reveal that MLLMs often generate incorrect answers even when they understand the visual content. To this end, we manually construct a benchmark with 12 categories and design evaluation metrics that assess the degree of error in MLLM responses even when the visual content is seemingly understood. Based on this benchmark, we test 15 leading MLLMs and analyze the distribution of attention maps and logits of some MLLMs. Our investigation identifies two primary issues: 1) most instruction tuning datasets predominantly feature questions that “directly” relate to the visual content, leading to a bias in MLLMs’ responses to other indirect questions, and 2) MLLMs’ attention to visual tokens is notably lower than to system and question tokens. We further observe that attention scores between questions and visual tokens as well as the model’s confidence in the answers are lower in response to misleading questions than to straightforward ones. To address the first challenge, we introduce a paired positive and negative data construction pipeline to diversify the dataset. For the second challenge, we propose to enhance the model’s focus on visual content during decoding by refining the text and visual prompt. For the text prompt, we propose a content guided refinement strategy that performs preliminary visual content analysis to generate structured information before answering the question. Additionally, we employ a visual attention refinement strategy that highlights question-relevant visual tokens to increase the model’s attention to visual content that aligns with the question. Extensive experiments demonstrate that these challenges can be significantly mitigated with our proposed dataset and techniques. The benchmark, training set, and code will be available.
Abstract
Due to the space constraints of the main manuscript, this Supplementary Material provides a comprehensive presentation of additional details and analyses. The content is structured as follows: (1) a detailed overview of the MMVU test set, including its construction criteria and a preliminary evaluation conducted on a subset of the test set; (2) findings from initial exploratory experiments to assess the test set’s effectiveness using a subset of the MMVU dataset; (3) an examination of the MLLM’s performance under various data combination strategies, prompt versions, and subsets of MMVU-Train of differing sizes, accompanied by an in-depth discussion on the construction of training prompts and their impact on model performance; (4) the training of an MLLM using the MMVU-Train dataset, validated on both a subset of the MMVU test set and other general datasets; (5) representative examples of paired positive and negative samples; and (6) the specific content of the ”Instruction” prompt employed during model testing.
Figure 1: (a) Examples illustrating MLLMs can accurately understand the visual content but provide incorrect responses. In each example, we show a pair of so-called positive and negative questions. The model can answer the positive questions (indicated as green), demonstrating that it understands the image, but fails to generate the correct answers on the negative question (indicated as red). This paper investigates this phenomenon. (b) Comparison of the model’s accuracy in response to positive and negative questions. (c) Performance comparison of MLLMs after fine-tuning with our dataset. The metric utilized here is the Response Accuracy (RA) in Sec. 3.3.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have garnered significant attention due to their impressive capabilities in addressing diverse multimodal tasks, particularly in visual comprehension and reasoning [72, 4, 26]. Despite these strengths, MLLMs sometimes produce responses that deviate from the image content, resulting in what is termed “hallucinations”. Previous studies [18] like MADBench [67], MME [21], POPE [47], ConBench [83], and NaturalBench [41] introduce datasets for evaluating the accuracy or consistency of MLLMs’ responses under image or text perturbations. However, they do not distinguish between the causes of incorrect responses. We observe that incorrect responses may not just stem from the MLLMs not understanding the visual content. Instead, we find that it is quite prevalent for an MLLM to understand the content of the image but yet still generate incorrect answers, as illustrated in Fig. 1 (a)). Our findings suggest that subtle phrasing in questions can introduce biases that hinder MLLMs from leveraging their understanding accurately. This raises two critical questions for our research: 1) Why does the model understand the visual content but provide incorrect answers? 2) How can we correct this error-prone behavior of MLLMs?
To explore this phenomenon, we introduce the MLLM Misresponses despite Visual Understanding (MMVU) benchmark, as shown in Fig. 3. Each sample is manually labeled with one of 12 types of questions, accompanied by paired positive and negative questions. Positive questions are designed to directly and objectively inquire about the observable content within an image, such as its types, attributes, poses, movements, or scenes, based on the actual and verifiable elements present, without introducing hypothetical alterations, irrelevant details, or erroneous premises. These questions aim to evaluate the model’s comprehension of the visual content. In contrast, negative questions involve hypothetical modifications, irrelevant descriptions, or incorrect assumptions about the image, which can lead to confusion or misinterpretation of the visual information. As a result, negative questions may induce incorrect responses, as they present scenarios that are not directly related to the image’s actual content. To ensure that the MLLMs select the most appropriate response based on their understanding of the visual content, we design multiple-choice questions with correct answers and carefully crafted distractors that include explanatory content, reducing the reliance on random or incorrect guesses. Additionally, MMVU comes with novel metrics to measure the likelihood of an MLLM’s failure when influenced by a negative question.
We evaluate 15 leading MLLMs on the MMVU benchmark, as shown in Tab. 1. We also analyze the attention maps and logits of some MLLMs when faced with different types of questions during the decoding process, as depicted in Fig. 4. Our findings reveal that most MLLMs respond unsatisfactorily to negative questions, demonstrating a biased response toward positive answers regardless of the question type. This phenomenon can be attributed to two main issues: 1) datasets in this area of research are predominantly composed of positive question-answer pairs, leading to significant bias in MLLMs toward positive questions due to the scarcity of negative samples, and 2) there is an unbalanced distribution of attention, with models prioritizing system and question tokens while significantly neglecting visual tokens. We further observe that attention scores between question and visual tokens as well as the model’s confidence in the correctness of the answers are lower in response to negative than positive questions.
To address the first issue, we introduce a data construction pipeline that extracts visual information from images, such as text, numbers, objects, attributes, relationships, and contextual elements at local and global levels. Based on this information, we construct 112k pairs of positive and negative samples to form the MMVU-Train dataset. For the second issue, we aim to enhance the model’s attention to visual content during decoding by implementing two key strategies: 1) Content Guided Refinement (CGR). We employ a two-step reasoning process where the model extracts detailed information from the image and then formulates responses based on this analysis alongside the visual content. 2) Visual Attention Refinement (VAR). To guide the MLLM in focusing on critical visual features relevant to the question, we extract attention scores and apply a mask to help focus on significant areas while masking the less important ones.
In summary, our contributions are as follows: 1) We identify and analyze the issue of MLLMs generating incorrect answers despite evidence of visual comprehension, attributing this to specific underlying biases. We introduce the MMVU benchmark and new metrics to evaluate MLLMs’ comprehension and resilience to negative questions. 2) We introduce a data construction pipeline in MMVU that generates pairs of positive and negative samples for training to enhance the model’s robustness to negative questions. Furthermore, we propose CGR and VAR strategies to optimize the model’s focus on relevant visual content and enhance response accuracy. 3) Extensive experiments reveal that our dataset and strategies effectively mitigate the susceptibility of MLLMs to negative questions.
@@ 2 Related Works
Figure 2: The MMVU dataset consists of a benchmarking dataset for evaluating models as well as a training dataset. The former is curated by human annotators together with the appropriate metrics and analysis on the MLLM’s attention and logit behavior. Based on the experiments, we propose a data construction pipeline to build a training dataset and prompting strategies to enhance the accuracy of MLLM responses.
@@
Multimodal Large Language Models (MLLMs). Building on the success of LLMs, recent research has increasingly focused on MLLMs [2, 72]. This shift aims to achieve better cross-modality understanding and reasoning. Currently, a significant amount of research is exploring various aspects of MLLMs, including the design of structures [55, 48], data construction [37, 49, 85], training strategies [62, 33, 58], and lightweight MLLMs [26, 88]. For a comprehensive review of these topics, readers are encouraged to investigate surveys [82, 32, 7] dedicated to several areas.
@@
Evaluations and Analysis of MLLMs. The assessment of MLLMs [45] is mainly divided into two main categories: general capability [59, 87] and hallucination benchmarks [8]. General capability benchmarks can be divided into single-task and comprehensive benchmarks. Single-task benchmarks, such as VQA [23], VQA-v2 [23], GQA [29], and MS-COCO [71], are designed to assess the performance of models on a specific defined task. Comprehensive benchmarks, such as MMBench [57], SEED [42], MMMU [80], and MMStar [12] aim to assess MLLMs’ general multi-modal perception and reasoning capabilities. Hallucination benchmarks mainly focus on evaluating the discrepancy between generated responses and visual content [8, 47]. Most hallucination benchmarks focus on object [34, 52, 74, 46, 66, 67, 56, 31, 68, 24], knowledge [52, 56, 31, 24], relation [79, 13, 31, 68], attribute [79, 28, 67, 56, 31, 68, 24], and spurious images [25, 70]. Current benchmarks assume that correct answers reflect understanding, while incorrect ones indicate a lack of understanding. However, our findings show that this incorrect stems not only from visual misunderstanding but also from a lack of robustness to negative questions.
3 Methods
@@
Figure 3: Examples in the MMVU benchmark. POS and NEG denote the positive and negative questions, ANS denotes the answer.
3.1 An Overview of MLLMs
The architecture of most MLLMs consists of three primary components: the visual encoder, the connector, and the LLM. Given a system prompt $P$, an input image $I$, and a question $Q$, the process commences with a text tokenizer that converts the system prompt and questions into text tokens $T_{P}$ and $T_{Q}$, respectively. The image $I$ is then encoded into visual tokens $T_{V}$ by the visual encoder, which captures essential visual features. Subsequently, the connector integrates the visual tokens $T_{V}$ with the text tokens to align the multi-modal data. Finally, the concatenated tokens (comprising text and visual tokens) are fed into the LLM to generate a response.
A significant challenge faced by current MLLMs is their ability to understand visual content while often providing incorrect responses to related questions. This paper aims to analyze the underlying causes of this issue and propose effective solutions. To facilitate this investigation, we introduce a MLLM Misresponses despite Visual Understanding (MMVU) benchmark, consisting of both positive and negative samples, as shown in Fig. 3. Each image in the benchmark is paired with two questions: a positive question that assesses the MLLM’s comprehension of the visual content and a negative question intended to confuse the model. This benchmark allows for a comprehensive analysis of the MLLMs’ performance in accurately answering questions and their vulnerability to misleading prompts. We establish two key metrics to evaluate the model’s ability to respond correctly to both positive and negative questions and to assess its performance when confronted with misleading queries despite its understanding of the visual content. Additionally, we analyze the attention mechanism and logit distribution of some open-source MLLMs to gain deeper insights into why this phenomenon happens.
@@ 3.1 An Overview of MLLMs
Table 1: Comparison to state-of-the-art MLLMs on the MMVU test set. Abbreviations: Char/Num. (Character/Number), Pres. (Presence), Color/Tex. (Color/Texture), Num. (Number), Shape (Shape), Posture (Posture), Pos. (Position), Abstract. (Abstract Knowledge), Concrete. (Concrete Knowledge), Expert (Expertise), Act. (Activity), Rel. (Relationships). $\uparrow$: higher is better, $\downarrow$: lower is better. Bolding and underlining indicate the best and second-best performance, respectively.
@@
GPT4o [65] [Uncaptioned image] 76.83 74.70 67.74 38.27 76.39 78.57 49.30 72.06 70.31 53.85 74.44 43.66 65.06
Qwen-VL-max-0809 [6] 67.07 60.87 61.29 34.48 76.00 62.07 42.86 69.57 72.88 67.74 70.73 54.55 63.36
Claude3.5-Sonnet-0620 [5] 64.63 53.01 66.13 38.27 56.94 67.86 38.03 64.71 64.06 44.62 67.78 35.21 55.32
Gemini-1.5-Pro [72] 47.56 36.14 41.94 23.46 43.06 40.48 15.49 60.29 56.25 32.31 50.00 28.17 39.53
@@
Ovis1.6-Gemma2-9B [60] [Uncaptioned image] 71.95 74.70 69.35 58.02 65.28 76.19 57.75 77.94 68.75 47.69 75.56 52.11 66.74
Llama3.2-90B-Vision-Instruct [20] [Uncaptioned image] 60.98 80.72 74.19 55.56 70.83 79.76 60.56 67.65 76.56 27.69 82.22 56.34 66.74
MiniCPM-2.6V [76] 69.51 72.29 69.35 48.15 51.39 73.81 45.07 70.59 71.88 38.46 80.00 35.21 61.14
LLaVA-OneVision-7B [43] 62.20 67.47 70.97 41.98 58.33 70.24 49.30 77.94 71.88 35.38 75.56 42.25 60.58
Idefics3-8B-Llama3 [38] 57.32 61.45 61.29 43.21 69.44 64.29 40.85 61.76 64.06 30.77 68.89 36.62 55.43
Pixtral-12B-2409 [3] 48.78 65.06 66.13 49.38 50.00 54.76 56.34 61.76 65.63 24.62 66.67 39.44 54.31
Phi3.5-Vision-Instruct [1] 54.88 53.01 61.29 37.04 58.33 55.95 35.21 60.29 64.06 29.23 68.89 38.03 51.62
InternVL2-8B [14] 46.34 65.06 53.23 33.33 56.94 53.57 39.44 61.76 65.63 29.23 66.67 21.13 49.72
LLaVA-v1.6-Vicuna-7B [54] 39.02 56.63 59.68 34.57 54.17 58.33 22.54 55.88 60.94 24.62 58.89 38.03 47.14
Cambrian-8B [73] 35.37 60.24 41.94 40.74 41.67 57.14 39.44 44.12 42.19 24.62 51.11 28.17 42.89
VILA1.5-13B [50] 14.63 30.12 35.48 12.35 27.78 42.86 8.45 29.41 29.69 16.92 43.33 14.08 25.76
@@
GPT4o [65] [Uncaptioned image] 8.69 16.22 16.00 38.00 11.29 12.00 32.69 18.33 19.64 23.91 16.25 35.42 19.53
Qwen-VL-max-0809 [6] 25.68 30.00 13.64 33.33 9.52 30.77 40.00 30.43 21.82 25.00 17.14 40.00 25.35
Claude3.5-Sonnet-0620 [5] 15.87 29.03 18.00 34.04 22.64 21.92 34.15 18.52 19.61 25.64 18.67 19.35 22.69
Gemini-1.5-Pro [72] 26.42 42.31 38.10 53.66 31.11 37.04 56.00 25.45 26.53 34.38 31.82 33.33 35.11